OpenStudy (ajprincess):

Note to the viewers: This is nt a question. I made a mistake in certain portion of my previous post which was a tutorial on trigonometric identities and this is the correct version of it.

OpenStudy (ajprincess):

jst a moment then.

OpenStudy (ajprincess):

\[sin(A+B)=sinAcosB+cosAsinB\] \[cos(A+B)=cosAcosB-sinAsinB\] \[sin(A-B)=sinAcosB-cosAsinB\] \[cos(A-B)=cosAcosB+sinAsinB\]

OpenStudy (valpey):

You can also write: \[sin(A\pm B)=sinAcosB\pm cosAsinB\] \[cos(A\pm B)=cosAcosB\mp sinAsinB\]